OpenStudy (anonymous):

find f'(a) f(x) = sqrt(2 − 2x) This is calculus. I'll give a medal please! show your work please so I can understand

OpenStudy (anonymous):

The derivative of a sum is the sum of the derivatives. The derivative of a constant is 0. The derivative of a function times a constant multiple is that multiple times the derivative of the function. The derivative of x^n w.r.t. x is nx^(n-1). These are the properties you should use.

OpenStudy (anonymous):

of course!

OpenStudy (anonymous):

so you could begin by using the difference quotient, correct?

OpenStudy (anonymous):

how did you derive f prime of x here?

OpenStudy (anonymous):

hm, but I am confused.. I see you wrote f(x) in exponential form, but I don't know why you have multiplied it by 1/2 and by -2? I must've missed something in your steps.

OpenStudy (anonymous):

Power rule and chain rule. d/du u^(1/2) = 1/2 u^(-1/2) d/dx (2 - 2x) = -2

OpenStudy (anonymous):

But my class hasn't gotten to this topic yet. Is there any other why to solve it?

OpenStudy (anonymous):

Um, rude. Just seeking help, most we've done is find tangent lines for given points.